# FlightBERT++: A Non-autoregressive Multi-Horizon Flight Trajectory Prediction Framework

### v2

@@ Conclusion
In this paper, we present a novel FlightBERT++ framework to perform the multi-horizon FTP task in a non-autoregressive manner. The proposed framework not only inherits the superior representation ability of the BE in the FlightBERT framework but also develops an innovative multi-horizon FTP model. Benefiting from the proposed differential prediction paradigm, the FlightBERT++ can mitigate the high-bit errors of the BE representation and achieve a significant reduction in outliers than that of the FlightBERT. Furthermore, FlightBERT++ exhibits the outstanding capability of trajectory representation through the visualizations of the trajectory embeddings, which provides interpretability for its superior performance. In addition, the proposed differential prompt mechanism is also confirmed to contribute to the performance improvements of the FTP task. In summary, FlightBERT++ achieves significant performance improvements and outperforms the comparison baselines across all evaluation metrics, especially in multi-horizon prediction scenarios.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (NSFC) under grants No. 62371323, U20A20161, U2333209, and 62001315.

### v4

@@ VI Conclusion and Future Work
In this paper, we present a novel FTP framework, called FlightBERT++, to perform the multi-horizon trajectory prediction in a non-autoregressive manner. The proposed framework not only inherits the superior representation ability of the BE in the FlightBERT framework but also develops an innovative multi-horizon neural network architecture for FTP tasks. The FlightBERT++ framework is implemented by a generalized encoder-decoder architecture, in which an additional HACG is designed to generate the multi-horizon contexts by considering the prior horizon information. Benefiting from the proposed GC representation and differential prediction paradigm, the FlightBERT++ can mitigate the high-bit misclassification of the BE representation and achieve a significant reduction in outliers than that of the FlightBERT. Furthermore, FlightBERT++ exhibits the outstanding capability of trajectory representation through the visualizations of the trajectory embeddings, which provides interpretability for its superior performance. In addition, the proposed differential prompt mechanism is also confirmed to contribute to the performance improvements of the FTP task. In summary, the proposed framework achieves significant performance improvements and outperforms the comparison baselines across most evaluation metrics, especially in multi-horizon prediction scenarios.
In the future, we will further focus on technical improvements of the GC representation and the FTP tasks. In addition, the practical applications of the FTP in the ATC domain are also interesting research topics, such as conflict detection, and traffic flow prediction.